# LASLiN: A Learning-Augmented Peer-to-Peer Network

### v1

@@ Abstract
We introduce a learning-augmented peer-to-peer (P2P) network design that leverages the predictions of traffic patterns to optimize the network’s topology. While keeping formal guarantees on the standard P2P metrics (routing path length, maximum degree), we optimize the network in a demand-aware manner and minimize the path lengths weighted by the peer-to-peer communication demands. Our protocol is learning-augmented, meaning that each node receives an individual, possibly inaccurate prediction about the future traffic patterns, with the goal of improving the network’s performances. In this paper, we strike a trade-off between significantly improved performances when the predictions are correct (consistency) and polylogarithmic performances when the predictions are arbitrary (robustness). Our protocol builds upon a generalization of skip lists to networks.
We have two main contributions. First, we consider the centralized setting and show that the problem of constructing an optimum static skip list network (SLN) is solvable in polynomial time and can be computed via dynamic programming. This problem is the natural demand-aware extension of the optimal skip list problem; in particular, we use it to generate the predictions of our learning-augmented protocol.
Second, we introduce the Uniform P2P protocol which generalizes skip list networks (SLN) by allowing the nodes to have any real-valued height instead of discrete ones. We show that Uniform achieves state-of-the-art performances: logarithmic routing and maximum degree, both with high probability. We then use Uniform to build a learning-augmented P2P protocol in order to incorporate demand-awareness, leading to our main contribution, LASLiN. We prove that the performances of LASLiN are consistent with those of an optimum static SLN with correct predictions, and are at most a logarithmic factor off the state-of-the-art P2P protocols if the predictions are arbitrary wrong. For the special case of highly sparse demands, we show that LASLiN achieves an exponential improvement in case of correct predictions, and asymptotically matches the performances of state-of-theart, demand-oblivious P2P protocols in case of adversarial predictions.
keywords:
learning-augmented algorithms, peer-to-peer networks, skip list graphs
@@ 1 Introduction
Peer-to-peer (P2P) protocols are one of the key technological advancements that allowed decentralized and distributed systems to scale. A vast body of work has been devoted to optimize the efficiency of these protocols [19, 27, 3, 16, 17]. Traditionally, the efficiency of a P2P protocol is estimated with two key metrics which are the routing path length between a pair of nodes and the maximum degree of each node in the network. State-of-the-art P2P protocols achieve both a logarithmic routing path length and a logarithmic maximum degree [29, 13].
Up until recently, most work on P2P protocols focused on demand-oblivious worst-case scenarios, where no prior information about the possible future demand is known. However, a new line of work studies demand-aware networks [28, 5, 12] which can adapt their connectivity properties to better fit the current traffic. Noting that most networking traffics are sparse and partly predictable [4], the idea is to detect the few highly active nodes and move them topologically closer to reduce the overall delays. The model includes a demand matrix which indicates the amount of traffic between any two nodes in the network; A goal of a demand-aware protocol is then to minimize the total weighted path length [28].
This paper presents a P2P network design where each node can adapt its set of neighbours to better fit the future traffic demands. We do not assume that the demand matrix is known by the nodes; Rather, we stick with the distributed nature of P2P protocols and assume that each node has access to a local prediction about the future demands — the prediction may only consist of partial information about the future traffic, an integer number in the case of this work. In our model, a node may obtain a new prediction at any time to use at its own discretion, independently from the other nodes. Crucially, the predictions are untrusted and we evaluate our P2P protocol in the learning-augmented setting with two metrics: consistency and robustness [20]. The consistency of the algorithm is its performance when the predictions are always perfect, its robustness is its performance when the predictions are worst possible.
@@ 1.1 Our Setup
P2P Networks
A P2P protocol is an algorithm distributed over multiple nodes that provides routing facilities between its member nodes. On top of a physical, static network, the P2P protocol keeps track of local routing information stored inside each node.
This paper studies P2P network protocols in which each node $u$ holds a height variable $h_{u}$. In addition, we assume that each node holds a unique identifier (id) which defines a total order at any time: for any two different nodes $u$ and $v$, either $u<v$ or $v<u$. A Skip List Network (SLN[21] is a P2P design based on the well-known skip list data structure [25] which establishes the routing informations solely based on the heights and ids of the nodes. In a SLN, two nodes $u$ and $v$ form an edge, i.e., store each other’s id, if and only if there is no node $x$ such that $u<x<v$ and $h_{x}\geq\min(h_{u},h_{v})$ — see Figure 1.
P2P Networks Constraints
Following the overall consensus [13, 23] on P2P networks, we restrict our focus to P2P protocols with the following properties: (1) any node should be able to join or leave the network at any time, (2) the routing should be local, i.e., each new hop through the network is decided solely by one node, and (3) the routing should be oblivious, i.e., the id of the destination node is sufficient to route the message.
Stated the above requirements, we want the network to still work at scale when the number of nodes in the network becomes large. Noting $n$ the number of nodes at the considered time, we aim at networks with:
  • small routing path lengths. Given any two nodes $u$ and $v$, at all time, the routing path length between $u$ and $v$ is ideally logarithmic in $n$ with high probability.
  • small degrees. At all time, the number of known nodes for any node, i.e., the maximum degree, is ideally logarithmic in $n$ with high probability.
All the P2P protocols described of this paper respect those two criteria. To evaluate the performances of our P2P protocols, we compare them to idealized P2P protocols with the restriction that they need to have polylogarithmic path length and maximum degree (see Definition 2.4).
Traffic Demand
@@
At a given time, we assume that the traffic demand between the nodes is modelled with a demand matrix $W\in M_{n}(\mathbb{R})$, where $W(u,v)$ is the frequency of communications from $u$ to $v$ [28].
Learning-Augmented Protocols
This paper introduces and studies learning-augmented P2P protocols. In the general case, a learning-augmented algorithm is an algorithm that takes an additional input, the prediction, to improve its performances. Crucially, the prediction is untrusted. This means that, with regard to the other inputs of the algorithm, the prediction may be perfect (all things being equal, no other prediction results in a better performance), adversarial, or anything in between.
In our design of learning-augmented P2P protocols, we assume that each node can obtain a prediction at any time, independently from one another. The origin of the prediction is not the focus of this paper; it may come from a centralized server handing out traffic predictions to the nodes, or it may be locally computed by each node based on its own traffic history. In this work, the input prediction is an integer interpreted as the height of the node in the SLN.
1.2 Contributions and Technical Novelty
@@
The contributions of this paper are threefold. First, we formulate the Optimum Static Skip List Network (OSSLN) problem and prove it can be solved in polynomial time, with an algorithm based on dynamic programming (Theorem 3.3). This result extends the previous work of [22] to the case with pairwise demands in line with recent demand-aware results [28]. Later on, this algorithm demonstrates that the task of finding optimal predictions for our learning-augmented P2P protocol is also polynomial.
012312345678
Figure 1: A representation of a Skip List Network (SLN) with eight nodes. The nodes have ids ranging from $1$ to $8$. The height of a rectangle equals the height of its corresponding node; Notice that all heights have integer values as this is an SLN. The double-arrowed, horizontal lines are the edges of the SLN.
0112345678
Figure 2: A continuous SLN with real-valued heights in $(0,1)$.
@@
Figure 3: A Continuous SLN obtained by adding node-by-node the heights of the SLN in Figure 1 and the heights of the Continuous SLN in Figure 2. This is a possible outcome of LASLiN where the predicted heights are those of Figure 1 and the uniform, random heights are those of Figure 2.
01212345678
Figure 4: A SLN with all equal heights, effectively turning the SLN into a simple chained list. This SLN has a linear routing path length. This worst-case performance happens when all the nodes get the same predicted height.
01212345678
Figure 5: The addition of continuous, random heights restores the good performances.
As our second contribution, we present a new P2P design called Continuous Skip List Network (Continuous SLN, Definition 4.1), which generalizes Skip List Networks [6] (SLN, Definition 2.1) and skip lists [25]. We show that a SLN with real-valued heights enables a functioning P2P protocol in that we implement the three basic primitives route, join and leave with state-of-the-art, logarithmic message complexity. We note that all known P2P designs that are based on skip lists [16, 14, 3] historically perceive it as chained lists stacked on each other and explicitly use that discrete nature in the protocol implementation, we show that this is not necessary. We then present a Continuous SLN protocol called Uniform that draws its heights uniformly at random in the interval $(0,1)$ and show that Uniform matches the performances of state-of-the-art P2P protocols with both routing path lengths and maximum degrees lower than $O(\log n)$ with high probability.
As our third contribution, we present a learning-augmented P2P protocol, LASLiN (Learning-Augmented Skip List Network), which combines our Uniform protocol with a predicted SLN to strike consistency and robustness guarantees (Definition 2.8). For each node, LASLiN takes as prediction an integer between $1$ and $\lceil\ln n\rceil$ and sets it as its integer heights at initialization time; it then draws a number in $(0,1)$ uniformly at random and adds it to the integer heights to obtain its final heights as a Continuous SLN. We show that LASLiN is comparable to an optimum static SLN when the predictions are correct ($O(1)$-consistent) and at worst a $\ln n$ factor off a demand-oblivious P2P protocol with arbitrary predictions ($O(\log^{2}n)$-robust). Finally we show that, in the special case of a very sparse demand matrix (see Definition 6.5), the version of LASLiN with binary predictions always outperforms any demand-oblivious P2P protocol with high probability: LASLiN obtains an exponentially better cost with correct predictions while it asymptotically matches the cost of the state-of-the-art P2P protocols for arbitrary predictions (see Theorem 6.6). We provide simulations to assess the performance of LASLiN on Zipfian demand matrices.
@@ 2 Model and Metrics
In this section, we formally define our model and the metrics we use to evaluate our P2P protocols.
Definition 2.1 (SLN).
@@ A Skip List Network (SLN) $\mathcal{N}$ is a triple $(V,h,E)$ where:
  • $h$ is a height function $h:V\longrightarrow\mathbb{N}^{+}$;
  • $E=\{\{u,v\}\;|\;\forall x\in V\;s.t.\;u<x<v,\;h_{x}<\min(h_{u},h_{v})\}$.
Definition 2.2 (Neighbor and Degree).
We say that two nodes $u$ and $v$ are neighbors in a SLN $\mathcal{N}=(V,h,E)$ if $\{u,v\}\in E$. The degree of a node $u$ is its number of neighbors.
We redefine the greedy, local routing path proper to SLN.
Definition 2.3 (Routing Path).
Let $\mathcal{N}=(V,h,E)$ be a SLN and $u,v\in V$ be two different nodes. The routing path from $u$ to $v$ in $\mathcal{N}$ is the sequence of edges of $E$ $s=e_{1},e_{2}\dots e_{k}$ such that: (1) $u\in e_{1}$, (2) $v\in e_{k}$, and (3) $\forall i\in[k]$, if $u<v$ then $e_{i}=\{x,y\}$ where $y$ is the largest node neighbor of $x$ such that $x<y\leq v$, otherwise if $v<u$ then $e_{i}=\{x,y\}$ where $y$ is the smallest node neighbor of $x$ such that $v\leq y<x$.
More intuitively, the routing path goes to the next neighbor that is the closest to the destination (in terms of ids) without “flying over” the destination.
Definition 2.4 (Ideal P2P Protocol).
We say that a (potentially randomized) P2P protocol is Ideal if, at all times, the routing path length between two nodes and the maximum degree are polylogarithmic on $n$ with high probability, where $n$ is the number of nodes at the considered time. A quantity is polylogarithmic when there exists positive constants $c$ and $d$ such that the quantity is smaller than $d\cdot\log^{c}n$.
Definition 2.5 (Cost).
@@
Definition 2.6 (Optimum Static Skip List Network (OSSLN) problem).
@@
Given a set of node ids $V$ and a demand matrix $W$, the objective of the Optimum Static Skip List Network (OSSLN) problem is to find a height assignment $h:V\longrightarrow\mathbb{N}^{+}$ that minimizes the cost of the SLN $\mathcal{N}=(V,h,E)$.
Definition 2.7 (Learning-augmented P2P protocol).
A learning-augmented P2P protocol ALG is a P2P protocol whose initialization routine takes an additional parameter $p$, the prediction.
In our model, a node that wishes to update its prediction as it already joined the network simply leaves, re-intializes and joins again.
Definition 2.8 (Consistency and Robustness).
We say that a learning-augmented SLN protocol is
@@
  • $\alpha$-consistent if, for all demand matrices, there exists an assignment of predictions to the nodes such that the expected cost of the protocol is at most $\alpha$ times the cost of an ideal optimum static SLN.
@@
  • $\beta$-robust if, for all demand matrices and for all assignment of predictions, the expected cost of the protocol is at most $\beta$ times the cost of an ideal optimum static SLN.
3 Optimum Static Skip List Network (OSSLN)
@@
Given a set of node ids $V$ and a demand matrix $W$, the objective of the Optimum Static Skip List Network (OSSLN) problem is to find a height assignment $h:V\longrightarrow\mathbb{N}^{+}$ that minimizes the total communication cost (as defined in (2.5)) of the SLN $\mathcal{N}=(V,h,E)$ (as defined in (2.1)):
$OSSLN(V,W)=\min_{h}\;Cost(\mathcal{N},W).$ (1)
@@
In [28], the authors presented an algorithm based on dynamic programming to find the Optimum static distributed Binary Search Tree (OBST) in polynomial time ($O(n^{3})$). This algorithm leverages a crucial property: once a vertex $v$ is chosen as the root of a tree induced by the vertices in a given interval, the subproblems induced by the vertex id intervals to the left and to the right of $v$ can be solved independently. The OSSLN problem exhibits a similar structure, but with crucial difference.
@@
If we define a pivot vertex $x\in[l,r]$, $1\leq l\leq r\leq n,$ such that $h_{x}=\max_{w\in[l,r]}h_{w}$, $x$ also partitions the OSSLN problem into two subproblems (on the intervals $[l,v)$ and $(v,r]$). However, a critical distinction arises when $h_{x}<\max_{w\in[l,r]}h_{w}$, in which case $x$ will not be a part of all the paths between vertex pairs $(u,v)$ $u\in[l,x),v\in(x,r]$. Specifically, for a vertex $x\in[l+1,r-1]$ such that $h_{l},h_{r}>h_{x}$, any path between vertices $(u,v)$ where $u\leq l$ and $v\geq r$ does not include $x$. This changes the optimal substructure of the problem, requiring a more sophisticated dynamic programming approach.
@@
Given a SLN $\mathcal{N}$, we define a path $P_{h}(l,r)$ to be the set of vertices in the routing path between the vertices $l$ and $r$ (as defined in (2.3)), excluding $l$, i.e., $l\notin P_{h}(l,r)$. This is a crucial simplification that allows the problem to be decomposed. Under this assumption, we have $d(l,r)=|P_{h}(l,r)|$. W.l.o.g., we assume that $P_{h}(l,r)=P_{h}(r,l)$.
@@
Next, we define $L(x,h)=\min\{i\leq x\;|\;h_{j}\leq h_{x},\;\forall j\in[i,x]\}$, i.e., the vertex of minimum id between the set of all (consecutive) vertices of height lower or equal to $h_{x}$, to the left of $x$, and, analogously, $R(x,h)=\max\{i\geq x\;|\;h_{j}\leq h_{x},\;\forall j\in[x,i]\}$, i.e., the vertex with maximum id between the set of all vertices of height lower or equal to $h_{x}$, to the right of $x$. A vertex $x\in V$ belongs to a path $P_{h}(u,v)$ if and only if there does not exist any pair of vertices $(u^{\prime},v^{\prime})$ such that $u\leq u^{\prime}<x$, $x<v^{\prime}\leq v$, and (both) $h_{u^{\prime}},h_{v^{\prime}}>h_{x}$.
@@
$P_{x}(h)$ $=$ $\{(u,v)\mid u<x\;\wedge\;v\in[x,R(x,h)]\}\;\cup$
$\{(u,v)\mid u\in[L(x,h),x)]\;\wedge\;v>R(x,h)\}$
Proof 3.2.
If $u<x$ and $x\leq v\leq R(x,h)$, then, by definition of $R(x,h)$, $h_{x}\geq h_{v}\geq\min(h_{u},h_{v})$. Since $x$ lies between $u$ and $v$ and there are no edges “above" $h_{x}$ in the interval $[u,v]$ (i.e., $\nexists(w,z)\in E$, s.t. $h_{w}>h_{x}$ and $h_{z}>h_{x}$, $w,z]\subseteq[u,v]$), $x$ must be a part of $P_{h}(u,v)$.
Similarly, if $L(x,h)\leq u<x$ and $v>R(x,h)$, then, again by definition of $L(x,h)$, $h_{x}\geq h_{u}\geq\min(h_{u},h_{v})$. Since $x$ lies between $u$ and $v$ and there are no edges "above" $h_{x}$ in the interval $[u,v]$, $x$ must be a part of $P_{h}(u,v)$
@@
Because $P_{h}(u,v)$ is the shortest path from $u$ to $v$, it follows that $x\notin P_{h}(u,v)$.
@@
$\text{Cost}(\mathcal{N},W)$ $=$ $\sum_{u,v\in V}W(u,v)\cdot d(u,v)$
$=\sum_{u,v\in V}\sum_{x\in P_{h}(u,v)}W(u,v)$ $=$ $\sum_{x\in V}\sum_{(u,v)\in P_{x}(h)}W(u,v).$
This means that the total cost is equal to the sum of contributions of all vertices $x\in V$, where the contribution of a vertex $x$ to the total cost is defined as:

### v2

# A Learning-Augmented Overlay Network

@@ Abstract
This paper studies the integration of machine-learned advice in overlay networks in order to adapt their topology to the incoming demand. Such demand-aware systems have recently received much attention, for example in the context of data structures (Fu et al. in ICLR 2025, Zeynali et al. in ICML 2024). We in this paper extend this vision to overlay networks where requests are not to individual keys in a data structure but occur between communication pairs, and where algorithms have to be distributed. In this setting, we present an algorithm that adapts the topology (and the routing paths) of the overlay network to minimize the hop distance travelled by bit, that is, distance times demand. In a distributed manner, each node receives an (untrusted) prediction of the future demand to help him choose its set of neighbors and its forwarding table.
This paper focuses on optimizing the well-known skip list networks (SLNs) for their simplicity. We start by introducing continuous skip list networks (C-SLNs) which are a generalization of SLNs specifically designed to tolerate predictive errors. We then present our learning-augmented algorithm, called LASLiN, and prove that its performance is (i) similar to the best possible SLN in case of good predictions ($O(1)$-consistency) and (ii) at most a logarithmic factor away from a standard overlay network in case of arbitrarily wrong predictions ($O(\log^{2}n)$-robustness, where $n$ is the number of nodes in the network). Finally, we demonstrate the resilience of LASLiN against predictive errors (ie, its smoothness) using various error types on both synthetic and real demands.
@@ 1 Introduction
Overlay networks are one of the key technological advancements that allowed decentralized and distributed systems to scale. A vast body of work has been devoted to optimize the efficiency of these networks Lua et al. (2005); Ratnasamy et al. (2001); Aspnes and Shah (2003); Harvey et al. (2003); Kiffer et al. (2021). Traditionally, the efficiency of an overlay network is estimated with two key metrics which are the routing path length between a pair of nodes and the maximum degree of each node in the network. State-of-the-art overlay networks achieve both a logarithmic routing path length and a logarithmic maximum degree Stoica et al. (2003); Fraigniaud and Gauron (2006).
Up until recently, most work on overlay networks focused on demand-oblivious worst-case scenarios, where no prior information about the possible future demand is known. However, noting that most networking traffics are sparse and partly predictable Avin et al. (2020a), a new line of work studies demand-aware networks Schmid et al. (2016); Avin et al. (2020b); Figiel et al. (2025) which can adapt their connectivity properties to better fit the current traffic.
On another front, there has been a recent surge in learning-augmented systems which, compared to traditional ones, take an additional parameter (the prediction) to improve their performances, with most recent examples in data structures Fu et al. (2025); Zeynali et al. (2024). There, the goal is to use a predicted demand for each node and adapt the system to better fit the incoming real demand. Since predicted and real demands may mismatch, they analyze what happens when the prediction is perfect (consistency), adversarial (robustness) and anything in between (smoothness, see Lykouris and Vassilvitskii (2018)).
This work considers an overlay network with $n$ nodes and assume it faces a certain real demand. Like previous works on learning-augmented systems, our goal is to use a predicted demand to answer the overall demand with delays (cost) smaller than classic, prediction-oblivious systems. We use the framework of consistency and robustness to assess the performances of our overlay algorithm; for that we compare the performances of our learning-augmented algorithm with those of a Skip List Network (SLN, 1) whose heights are optimal with respect to the real demand.
Unlike previous works however, our framework shows two crucial differences compared to other learning-augmented systems: our algorithm must (1) obey the standard constraints of overlay networks (Subsection 1.1), and (2) optimize on a pairwise demand representing the intensity of the exchanges between any two nodes (Subsection 1.1).
@@ 1.1 Our Setup
Network and Demand.
We model an overlay network by an undirected graph with $n$ nodes. An edge of the graph means that the two end-nodes can communicate directly without the help of other nodes, ie, they hold each other’s ids. This graph is an abstract representation of the routing possibilities between the nodes and may be different from the underlying, physical network that carries the communications.
@@
At the considered time, we model a demand between the nodes with a demand matrix $W\in M_{n\times n}(\mathbb{R})$, where $W(u,v)$ is the amount of communications from $u$ to $v$.
Overlay Network Algorithms.
An overlay network algorithm is an algorithm distributed over the nodes of the network that defines the connectivity between overlay nodes and how routing can be performed. On top of the physical, static network, the overlay algorithm keeps track of local routing information stored inside each node.
Prediction Model.
The initialization procedure of our overlay algorithm (run right before the join procedure) takes a prediction as an additional argument (see also Aradhya and Scheideler (2025)). As a newcomer node joins the overlay network, the prediction influences its neighborhood in the network to (ideally) fit the incoming demand best. In the case of this paper, the prediction is the (integer) height of the node in the considered skip list network. Adaptation to a changing demand (and thus changing prediction) is performed by leaving and re-joining the network.
Consistency, Robustness, and Smoothness.
The predictions are untrusted and should not be followed blindly by the nodes. In this context, we evaluate the performances of our algorithm with the two most frequently studied metrics called consistency and robustness Lykouris and Vassilvitskii (2018) (7). In more details, we derive overlay algorithms that provably improve their performance in case of globally accurate predictions (consistency), but still present good performances (poly-logarithmic in $n$ with high probability) in case of arbitrary predictions, even adversarial ones (robustness). In our distributed context however, we observe that the consistency metric may be too weak and its assumption that all nodes have good predictions unrealistic. We therefore consider a third metric, smoothness, which indicates the performances along the spectrum from consistency to robustness; in particular, what happens when the prediction is good but with a slight noise. We provide extensive analysis of the smoothness of our algorithm tested on synthetic and real demands and with different types of errors (Section 6).
1.2 Contributions
@@
The contributions of this paper are threefold. First, we formulate the Optimum Static Skip List Network (OSSLN) problem and prove it can be solved in polynomial time, with an algorithm based on dynamic programming (Theorem 2). This result extends the previous work of Martinez and Roura (1995) to the case with pairwise demands in line with recent demand-aware results Schmid et al. (2016). This algorithm demonstrates that the task of finding optimal predictions for our learning-augmented overlay algorithm is polynomial, which allowed us to run our evaluations.
As our second contribution, we present a new overlay design called Continuous Skip List Network (C-SLN, 9), which generalizes Skip List Networks Avin et al. (2020c) (SLN, 1) and the classic skip lists Pugh (1990). We show that a C-SLN offers a functioning overlay algorithm in that we implement the three basic primitives join, leave and route. We then present a specific (without prediction) C-SLN called Uniform that draws its heights uniformly at random in the interval $[0,1]$. We show that Uniform matches the performances of state-of-the-art overlay algorithms with both routing path lengths and maximum degrees in $O(\log n)$ with high probability.
As our third contribution, we present a learning-augmented overlay algorithm, LASLiN (Learning-Augmented Skip List Network), which combines our Uniform algorithm with a predicted SLN to strike consistency and robustness guarantees (7). For each node, LASLiN takes as prediction an integer between $1$ and $\lceil\ln n\rceil$ and sets it as its integer heights at initialization time; it then draws a number in $[0,1]$ uniformly at random and adds it to the integer heights to obtain its final heights as a C-SLN; it may then join the network using the join() procedure. We show that LASLiN is comparable to an optimum static SLN when the predictions are correct ($O(1)$-consistent) and at worst a $O(\log^{2}n)$ factor away from the optimum static SLN in case of adversarial predictions ($O(\log^{2}n)$-robust).
@@
Figure 1: A SLN with eight nodes. The node ids range from $1$ to $8$. The double-arrowed, horizontal lines are the edges of the SLN.
0123412345678
Figure 2: A C-SLN obtained by adding each height of Figure 2 with a random number in $[0,1]$. This is a possible outcome of LASLiN with predicted heights in green.
@@ 2 Model and Metrics
In this section, we formally define our model and the metrics we use to evaluate our overlay algorithms.
Definition 1 (SLN).
@@ A Skip List Network (SLN) $\mathcal{N}$ is a triple $(V,h,E)$ where:
  • $h$ is a height function $h:V\rightarrow\mathbb{N}^{+}$;
  • $E$ is the set of edges such that $\{u,v\}\in E$ if and only if
    $\forall x\in V\;\text{ s.t. }\;u<x<v,\quad h_{x}<\min(h_{u},h_{v})$
We redefine hereafter the greedy, local routing path specific to SLN. Intuitively, the routing path goes to the next neighbor that is the closest to the destination (in terms of ids) without “flying over” the destination.
Definition 2 (Routing path).
[…]
Definition 3 (Cost).
@@
Definition 4 (Optimum Static Skip List Network (OSSLN) problem).
@@
Given a set of nodes $V$ and a demand matrix $W$, the objective of the Optimum Static SLN (OSSLN) problem is to find a height assignment $h:V\rightarrow\mathbb{N}^{+}$ that minimizes the cost of the SLN $\mathcal{N}=(V,h,E)$.
Definition 5 (Overlay Algorithm).
An overlay algorithm is a distributed algorithm running on each node of the network, which provides three procedures:
  • initialization: Executed before a node joins the network. In this paper, sets the height of each node in the SLN.
  • join: Executed to join the network. Takes as argument the id of a node that is already in the network.
  • leave: Executed to leave the network.
Definition 6 (Learning-Augmented Overlay Algorithm).
A learning-augmented overlay algorithm ALG is an overlay algorithm whose initialization procedure takes an additional parameter, the input prediction. In this paper, the input prediction consists of an estimated height assignment for each node.
Definition 7 (Consistency and Robustness).
We say that a learning-augmented C-SLN algorithm is
@@
  • $\alpha$-consistent if, for all demand matrices $W$ and all set of nodes $V$, there exists a set of predictions (assignment of heights) such that the expected cost of the algorithm is at most $\alpha$ times the cost of an optimal optimum static SLN for $V$ and $W$.
@@
  • $\beta$-robust if, for any demand matrix $W$, any set of nodes $V$, and any set of prediction (assignment of heights), the expected cost of the algorithm is at most $\beta$ times the cost of an optimal optimum static SLN for $V$ and $W$.
3 Optimum Static Skip List Network
@@
Given a set of nodes $V$ and a demand matrix $W$, the objective of the Optimum Static Skip List Network (OSSLN) problem is to find a height assignment $h:V\rightarrow\mathbb{N}^{+}$ that minimizes the total communication cost (3) of the SLN $\mathcal{N}=(V,h,E)$ (1):
$\textsc{OSSLN}(V,W)=\min_{h}\;Cost(\mathcal{N},W).$ (1)
where $\mathcal{N}$ is an SLN with node set $V$ and heights $h$.
@@
In Schmid et al. (2016), the authors presented a dynamic programming algorithm to find the Optimum static distributed Binary Search Tree (OBST) in polynomial time ($O(n^{3})$). This algorithm leverages a crucial property: once a node $v$ is chosen as the root of a tree induced by the nodes in a given interval, the subproblems induced by the node id intervals to the left and to the right of $v$ can be solved independently. The OSSLN problem exhibits a similar structure, but with a crucial difference.
@@
If we define a pivot node $x\in[l,r]$, $1\leq l\leq r\leq n,$ such that $h_{x}=\max_{w\in[l,r]}h_{w}$, $x$ also partitions the OSSLN problem into two subproblems (on the intervals $[l,v)$ and $(v,r]$). However, a critical distinction arises when $h_{x}<\max_{w\in[l,r]}h_{w}$, in which case $x$ will not be a part of all the paths between node pairs $(u,v)$ $u\in[l,x),v\in(x,r]$. Specifically, for a node $x\in[l+1,r-1]$ such that $h_{l},h_{r}>h_{x}$, any path between nodes $(u,v)$ where $u\leq l$ and $v\geq r$ does not include $x$. This changes the optimal substructure of the problem, requiring a more sophisticated dynamic programming approach.
@@
Given a SLN $\mathcal{N}$, we define a path $P_{h}(l,r)$ to be the set of nodes in the routing path between the nodes $l$ and $r$ (2), excluding $l$, i.e., $l\notin P_{h}(l,r)$. This is a crucial simplification that allows the problem to be decomposed. Under this assumption, we have $d(l,r)=|P_{h}(l,r)|$. W.l.o.g., we assume that $P_{h}(l,r)=P_{h}(r,l)$.
@@
Next, we define $L(x,h)=\min\{i\leq x\;|\;h_{j}\leq h_{x},\;\forall j\in[i,x]\}$, i.e., the node of minimum id between the set of all (consecutive) nodes of height lower or equal to $h_{x}$, to the left of $x$, and, analogously, $R(x,h)=\max\{i\geq x\;|\;h_{j}\leq h_{x},\;\forall j\in[x,i]\}$, i.e., the node with maximum id between the set of all nodes of height lower or equal to $h_{x}$, to the right of $x$. A node $x\in V$ belongs to a path $P_{h}(u,v)$ if and only if there does not exist any pair of nodes $(u^{\prime},v^{\prime})$ such that $u\leq u^{\prime}<x$, $x<v^{\prime}\leq v$, and (both) $h_{u^{\prime}},h_{v^{\prime}}>h_{x}$.
@@
$P_{x}(h)$ $=$ $\{(u,v)\mid(u<x)\wedge v\in[x,R(x,h)]\}\cup$
$\{(u,v)\mid u\in[L(x,h),x)]\wedge(v>R(x,h))\}.$
Proof.
If $u<x$ and $x\leq v\leq R(x,h)$, then, by definition of $R(x,h)$, $h_{x}\geq h_{v}\geq\min(h_{u},h_{v})$. Since $x$ lies between $u$ and $v$ and there are no edges “above” $h_{x}$ in the interval $[u,v]$ (i.e., $\not\exists(w,z)\in E$, s.t. $h_{w}>h_{x}$ and $h_{z}>h_{x}$, $w,z]\subseteq[u,v]$), $x$ must be a part of $P_{h}(u,v)$.
Similarly, if $L(x,h)\leq u<x$ and $v>R(x,h)$, then, again by definition of $L(x,h)$, $h_{x}\geq h_{u}\geq\min(h_{u},h_{v})$. Since $x$ lies between $u$ and $v$ and there are no edges ”above” $h_{x}$ in the interval $[u,v]$, $x$ must be a part of $P_{h}(u,v)$
@@
Because $P_{h}(u,v)$ is the shortest path from $u$ to $v$, it follows that $x\notin P_{h}(u,v)$. ∎
@@
$Cost(\mathcal{N},W)$ $=$ $\sum_{u,v\in V}W(u,v)\cdot d(u,v)$
$=$ $\sum_{u,v\in V}\sum_{x\in P_{h}(u,v)}W(u,v)$
$=$ $\sum_{x\in V}\sum_{(u,v)\in P_{x}(h)}W(u,v).$
This means that the total cost is equal to the sum of contributions of all nodes $x\in V$, where the contribution of a node $x$ to the total cost is defined as: